OpenStudy (anonymous):

Can someone please check my work for this problem? So I'm doing the math part of an online chemistry lab for my homeschooling and I'm usually not good at math when it comes to chemistry. Okay, make that ever. I'm generally not good at math in general. But I think I might be figuring this one out! There are more problems than this, but since I think I know what I'm doing now, would someone see if I'm doing this problem right so I know I can go ahead and solve the rest without worry? Hold on a sec so I can post it all

OpenStudy (anonymous):

First, the data I have for the lab is METAL: Iron Mass of Metal: 34.720 Volume of water in the calorimeter: 26ml Initial temperature of water in the calorimeter: 25.4°C Temperature of hot water and metal in hot water bath: 100.7°C Final temperature reached in the calorimeter: 33°C So with that data, here's the problem 1. Calculate the energy change (q) of the surroundings (water) using the enthalpy equation qwater = m × c × ΔT. And here is my work. I had to show all steps including what I just did in words. IRON Qwater=26*4.18J/(°C*g)*(33°C-25.4°C) Multiply 26 by 4.18 (mass of the metal by the SHC of water) Qwater=108.68J/(°C*g)*(7.6°C) Multiply 108.68 by 7.6 (joules by temperature) Qwater=825.968J/(°C*g) Please tell me if this is right really quickly! Thank you!

OpenStudy (anonymous):

Mass of the metal is 34.720 not 26ml (volume)

OpenStudy (anonymous):

Oh... Thank you! <3

OpenStudy (anonymous):

So it'd be 145.1296 in place of 108.68?

OpenStudy (goformit100):

@Dharmaputra_993 Come On.

OpenStudy (anonymous):

@XxDailyDreamxX yes

OpenStudy (anonymous):

Thank you :) So what I have now is Qwater=34.720 * 4.18J/(°C*g)*(33°C-25.4°C) Multiply 34.720 by 4.18 (mass of the metal by the SHC of water) and subtract 25.4 from 33. Qwater=145.1296J/(°C*g)*(7.6°C) Multiply 145.1296 by 7.6 (joules by temperature) Qwater=1102.98496/(°C*g) Is that right?

OpenStudy (anonymous):

If you are trying to find the change in energy from the initial water to the final, then yes.

OpenStudy (anonymous):

Thank you! Um... I have one more question. I'm trying to do the second part of the assignment, and I'm not sure how to write this equation out. Would you please do the first step by writing out the formula I'll solve please or tell me what goes where? 2. Using the formula qmetal = m × c × ΔT, calculate the specific heat of the metal. Use the data from your experiment for the metal in your calculation.

OpenStudy (anonymous):

It involves the same data used for the other problem, but no matter what I do to try putting it together it doesn't look like it makes sense.

OpenStudy (anonymous):

@nubeer? What should I do? (Sorry if I'm bothering you, I'm rather new to using openstudy and I've never tagged anyone. But thank you for your time!)

OpenStudy (anonymous):

@XxDailyDreamxX , The energy of the metal q is the same as that of the water. Substitute the mass of the metal for m, (You're trying to find c), Substitute the change in temperature of the metal for ΔT. (Remember that we're assuming that no heat energy is lost in the process, q=q).

OpenStudy (anonymous):

@dimensionx So would it be like this? c=1102.98496/(34.720* 7.6)

OpenStudy (anonymous):

Your ΔT is incorrect, find the ΔT for the change in temperature of the metal from the hot bath to the final temperature of the water (and the metal).

OpenStudy (anonymous):

c=1102.98496/(34.720* 67.7)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you! :)

OpenStudy (anonymous):

Glad I could help :)